OpenStudy (anonymous):

if h(x)=2g(x)-(1)/(g(x)), g(0)=2 and g'(0)=-1, find the exact value of h'(0). please help and show all steps!

OpenStudy (anonymous):

For starters I would rewrite your h(x) function \[h(x)=2g(x)-\frac{1}{g(x)}\] using an exponent \[h(x)=2g(x)-(g(x))^{-1}\]. Then, differentiate, using the chain rule on the second term. \[h'(x)=2g'(x)+(g(x))^{-2} \cdot g'(x)\] so... \[h'(x)=2g'(x)+\frac{g'(x)}{(g(x))^2}\]. From here, you can just substitute zero for \[x\] to get \[h'(0)=2g'(0)+\frac{g'(0)}{(g(0))^2}\]. Substituting the values from the original problem will give you the final answer.